Accelerator Update for June 8, 2009 - June 10, 2009
Monday June 8
The day shift began with Operations monitoring the Tevatron (TeV) store 7122 (91.22E30), the Recycler stash (199.0E10), and the Antiproton Source stack (25.3E10), with Pbar not stacking due to a power supply LCW leak, with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, NuMI, and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
A Booster power supply (GMPS #2) was brought online. This work held off beam from the Proton Source for 2 hours during the morning.
Operations terminated store 7122 at 12:59 PM. Operations prepared D0 for an access.
Operations disabled the MTest beam at 5:10 PM.
Operations established store 7126 at 6:45 PM with an average initial luminosity (AIL) of 274.08E30.
Tuesday June 9
The midnight shift began with Operations monitoring store 7126 (131.02E30) and the stash (128.8E10), with Pbar stacking (9.5E10), and with NuMI and MiniBooNE taking beam.
Operations established beam to MTest beam at 4 AM.
Booster personnel resumed the GMPS #2 work. Beam was off for 2 hours during the morning.
Operations terminated store 7126 at 5:02 PM.
Operations put the TeV into shot setup and established store 7127 at 7:02 PM with an AIL of 278.73E30.
Wednesday June 10
The midnight shift began with Operations monitoring store 7127 (134.44E30) and the stash (219.6E10), with Pbar stacking (18.8E10), and with NuMI and MiniBooNE taking beam.
Operations established beam to MTest at 4:05 AM.
Operations terminated store 7127 at 6 AM and then put the TeV into HEP shot setup.
Operations established store 7128 at 7:43 AM with an AIL of 289.54E30.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator will allow various machine studies in preparation for the coming shutdown.
Comments and Suggestions
|last modified 06/10/2009 email Fermilab|
|Security, Privacy, Legal|